# DiffAdapt: Difficulty-Adaptive Reasoning for Token-Efficient LLM Inference

### v2

@@ 1 Introduction
Large Language Models (LLMs) emerged as powerful tools for complex reasoning tasks, spanning mathematical problem-solving (Lewkowycz et al., 2022a; Pan et al., 2024), code generation (Chen et al., 2021), and logical deduction (Wei et al., 2022). A key ingredient in this success is intermediate reasoning steps, often referred to as a “chain of thought” (CoT), before producing a final answer (Yu et al., 2024; Li et al., 2025d; Liu et al., 2024). However, this capability comes at a significant computational cost. Models typically generate a lengthy and elaborate chain of thought for every problem, a process sometimes called test-time scaling (Muennighoff et al., 2025; Chu et al., 2025).
Always generating long traces is fundamentally inefficient. It squanders resources on simple problems, while not necessarily providing sufficient resources for truly complex tasks (Qu et al., 2025; Sui et al., 2025; Li et al., 2025b). In this work, we introduce a framework to bridge this gap through systematic empirical analysis and adaptive inference strategy design. We first discover a U-shaped entropy pattern: high entropy on easy problems despite high accuracy, low entropy on problems with medium difficulty, and high entropy on hard problems reflecting uncertainty. Counterintuitively, models show high uncertainty on simple problems despite achieving high accuracy.
This motivates us to design a different inference strategy per each of this three distinct difficulty regions. We develop three simple inference strategies, each equipped with different generation length (i.e., max token length), prompt and decoding hyperparamenters (e.g., sampling temperature). The prompt designed for “easy” questions encourages models to answer succintly without overthinking, while the prompt for high difficulty questions instructs model to think carefully. We first conduct oracle experiments with these three templates. When allowed to choose an optimal strategy per question, models achieve 50% token savings while improving the model accuracy by over 10%.
@@ 3 Characterizing Overthinking Phenomenon
Recent work has identified that reasoning LLMs exhibit “overthinking” behavior (Ma et al., 2025; Sui et al., 2025; Qu et al., 2025), where models generate exceedingly lengthy solution when they can arrive at correct solutions much more succintly. Building upon this observation, we analyze this phenomenon from an entropy perspective, revealing a counterintuitive pattern where models demonstrate high uncertainty.
@@ 3.2 Results
Figure 1 (a) illustrates this U-shaped entropy curve across different difficulty levels using the DeepSeek-R1-Distill-Qwen-7B model on the DeepMath-103K dataset (He et al., 2025). We observe consistent overthinking patterns across multiple model architectures (see Appendix D for additional models).
@@ 4 Oracle Experiments with Three Inference Strategies
Figure 2: Our three inference strategy configurations. ||Max|| refers to the predefined maximum token budget for generation. Full prompts and detailed configurations can be found in the Appendix B.
@@
Figure 3: Plots summarizing task performance (y-axis) and efficiency (x-axis) on Qwen3-4B model with various inference strategies. In all datasets, we find oracle strategy (i.e., the strategy that achieves the correct answer while incurring minimal tokens or the strategy incurring minimal tokens if no strategy leads to correct answer) significantly outperforms any uniform setting.
Inference Strategy Set
We design the inference strategies to address the challenges revealed by our overthinking analysis. For Easy problems, we use lower temperature and reduced tokens to prevent the model from exploring unnecessary solution paths. Normal problems receive standard temperature with full token budget to enable comprehensive reasoning in the optimal region where the model’s uncertainty appropriately matches task complexity. For Hard problems, we apply the lowest temperature with moderate token limit to encourage focused, deterministic reasoning while preventing resource waste on problems at the model’s capability limit. Table 3 summarizes the specific configurations for each reasoning strategy, more details can be found in the Appendix B.
Experimental Setting
To establish the theoretical upper bound of our approach, we conduct an oracle experiment using Qwen3-4B across eight reasoning benchmarks. We evaluate on five mathematical reasoning tasks: GSM8K (Cobbe et al., 2021), MATH500 (Lightman et al., 2023), AIME 2024 & 2025, and OlympiadBench (He et al., 2024), along with three out-of-domain benchmarks: Minerva (Lewkowycz et al., 2022b), GPQA (Rein et al., 2024), and MMLU-Pro (Wang et al., 2024). For each input, we generates outputs from each of the three inference strategies, Easy, Normal, and Hard. The Oracle strategy is determined through a two-step process: first, we identify all strategies that yield a correct answer, and from this set, we then select the one with the minimal token consumption. The parameter configurations are set as the same as the fixed strategies in Table 3, and the max token limit is 32K. More visualization details can be found in Appendix E.
@@ 5.1 Stage 1: Data Generation with Proxy Model Sampling
For each problem $x$, we prompt the model to generate complete reasoning steps and final results with a maximum length of 32K tokens. We then compute the model’s uncertainty using the same generation entropy calculation described in Section 3. We perform this process with 10 sampling iterations per problem to ensure robust entropy estimates. This generation entropy serves as a proxy for task complexity. Grounded in the U-shaped entropy pattern in Section 3 and the Pareto-optimal Oracle analysis (Appendix C.1), we adopt the following heuristic labeling rule: Normal if correctness $\geq\,\alpha$ and entropy $\leq\,\beta$; Hard if correctness $\gamma$; Easy otherwise (typically the low-difficulty anomaly with moderate correctness but unexpectedly high uncertainty). The thresholds $\alpha,\beta,\gamma$ are set per model using a simple heuristic informed by the observed entropy–correctness distributions; we optionally perform a light sanity check on a small validation split to ensure label stability. Per-model values are reported in Appendix C.2.
@@ 6.1 Experimental Setup
Experimental Design. To demonstrate the probe’s ability to perform difficulty-adaptive reasoning, we mix benchmarks of different difficulty levels. We present averaged results across benchmarks with varying domains, in-domain (GSM8K, MATH500, AIME24&25, OlympiadBench) and out-of-domain (Minerva, GPQA, MMLU-Pro) to show adaptive performance across the difficulty spectrum. Each experiment is run three times; we report the mean across runs. For evaluation, we follow reliability protocols from Ye et al. (2025) [^2] https://github.com/GAIR-NLP/LIMO,Chen et al. (2025) [^3] https://github.com/IAAR-Shanghai/xVerify. Max token limits: for each model–benchmark pair, we first allow a generous cap (e.g., 32K) to observe the model’s longest output and then set a nearby rounded value as the per-benchmark max_tokens used in plots; see Appendix Table 5 for the finalized values and Appendix C.4 for details.
@@ 6.2 Reasoning LLMs Results
Performance Analysis
DiffAdapt consistently outperforms fixed strategies across all model architectures and domains. On in-domain tasks (Figure 5(a)), Qwen3-4B shows the largest improvements, particularly at higher token budgets. DeepSeek-R1-Qwen-7B demonstrates stable gains, while DeepSeek-R1-Llama-8B exhibits significant benefits across model families. On out-of-domain evaluation (Figure 5(b)), the framework remains effective: DiffAdapt delivers consistent improvements across models and domains, with gains becoming more pronounced as the maximum token budget increases. By contrast, the training-free dynamic baseline DEER performs comparably to the strongest fixed strategy on in-domain datasets but still lags behind DiffAdapt; under distribution shift (Minerva, MMLU-Pro, GPQA), DEER exhibits limited generalization and larger performance degradation. These results confirm DiffAdapt’s robustness across different architectures, scales, and domains.
Key Findings
The results reveal distinct performance patterns for different strategies. Easy strategies achieve high performance with minimal tokens but plateau quickly. Normal and Hard strategies improve continuously with increased computational budgets. This validates averaging across benchmarks of varying difficulty, as it captures the distinct computational requirements of different problem complexities. DiffAdapt exploits these patterns through adaptive strategy selection, consistently outperforming fixed approaches across all token limits. Compared with DEER, DiffAdapt delivers larger and more stable gains across token budgets and domains, indicating that difficulty-aware strategy selection generalizes better than confidence-based early exit. Models with larger inter-strategy performance differences show greater DiffAdapt improvements, confirming that adaptive selection benefits from strategy diversity.
@@ 6.4 Computational Efficiency Analysis
Token cost
From Table 1, we observe substantial efficiency gains from DiffAdapt across models: Qwen3-4B achieves 22.4% token savings, while DS-R1-Qwen-7B and ThinkPrune-7B reduce usage by around 10%. In contrast, DEER increases token usage relative to Medium (e.g., 27.5-27.5% on Qwen3-4B and 53.3-53.3% on DS-R1-Qwen-7B). The reason is mechanistic: DEER operates under a fixed token budget and chooses continuation length primarily by confidence/probability, which frequently drives generations to the maximum token cap rather than allocating tokens adaptively by problem difficulty.
Latency
@@ Acknowledgements
We thank the members of NYU ML2 for helpful feedback. This work was supported in part by NYU HPC resources, services, and staff expertise.
@@
Appendix A Limitations and Future Work
@@
Probe input sufficiency. Our probe uses hidden states from the question prefill only, which may not fully capture difficulty signals. A promising direction is to incorporate generation-time latent information (e.g., step-wise entropy/logits, attention/KV statistics, or early decoding features) as additional inputs. Such intrusive signals, however, may interrupt standard generation pipelines and introduce extra latency; designing low-overhead hooks is left for future work.
Appendix B Complete Reasoning Strategy
@@
B.1 Easy Strategy Prompt
@@
B.2 Normal Strategy Prompt
@@
B.3 Hard Strategy Prompt
For problems at the capability limit, we focus on efficient resource utilization:
@@ B.3 Hard Strategy Prompt
  • Approach: Strategic method outline to prevent unproductive reasoning loops
B.4 Design Rationale
@@
  • Easy strategy discourages overthinking by emphasizing directness and verification rather than extensive exploration
  • Normal strategy encourages systematic reasoning with full computational resources for optimal performance
  • Hard strategy promotes efficient strategy selection to avoid getting stuck in computational dead ends
For the hyperparameter choices like temperature and max tokens, we run the grid search to find the best configuration for each strategy.

### v3

@@ 1 Introduction
Large Language Models (LLMs) emerged as powerful tools for complex reasoning tasks, spanning mathematical problem-solving (Lewkowycz et al., 2022a; Pan et al., 2024), code generation (Chen et al., 2021), and logical deduction (Wei et al., 2022; TANG et al., 2026; Tang et al., 2025b). A key ingredient in this success is intermediate reasoning steps, often referred to as a “chain of thought” (CoT), before producing a final answer (Yu et al., 2024; Li et al., 2025d; Liu et al., 2024; Kong et al., 2025). However, this capability comes at a significant computational cost. Models typically generate a lengthy and elaborate chain of thought for every problem, a process sometimes called test-time scaling (Muennighoff et al., 2025; Chu et al., 2025; Tang et al., 2025a).
Always generating long traces is fundamentally inefficient. It squanders resources on simple problems, while not necessarily providing sufficient resources for truly complex tasks (Qu et al., 2025; Sui et al., 2025; Li et al., 2025b; Liu et al., 2025d; c; Zhang et al., 2023; Dong et al., 2025). In this work, we introduce a framework to bridge this gap through systematic empirical analysis and adaptive inference strategy design. We first discover a U-shaped entropy pattern: high entropy on easy problems despite high accuracy, low entropy on problems with medium difficulty, and high entropy on hard problems reflecting uncertainty. Counterintuitively, models show high uncertainty on simple problems despite achieving high accuracy.
This motivates us to design a different inference strategy per each of this three distinct difficulty regions. We develop three simple inference strategies, each equipped with different generation length (i.e., max token length), prompt and decoding hyperparameters (e.g., sampling temperature). The prompt designed for “easy” questions encourages models to answer succinctly without overthinking, while the prompt for high difficulty questions instructs model to think carefully. We first conduct oracle experiments with these three templates. When allowed to choose an optimal strategy per question, models achieve 50% token savings while improving the model accuracy by over 10%.
@@ 3 Characterizing Overthinking Phenomenon
Recent work has identified that reasoning LLMs exhibit “overthinking” behavior (Ma et al., 2025; Sui et al., 2025; Qu et al., 2025), where models generate exceedingly lengthy solution when they can arrive at correct solutions much more succinctly. Building upon this observation, we analyze this phenomenon from an entropy perspective, revealing a counterintuitive pattern where models demonstrate high uncertainty.
@@ 3.2 Results
Figure 1 (a) illustrates this U-shaped entropy curve across different difficulty levels using the DeepSeek-R1-Distill-Qwen-7B model on the DeepMath-103K dataset (He et al., 2025). We observe consistent overthinking patterns across multiple model architectures (see Appendix F for additional models).
@@ 4 Oracle Experiments with Three Inference Strategies
Table 1: Our three inference strategy configurations. ||Max|| refers to the predefined maximum token budget for generation. Full prompts and detailed configurations can be found in the Appendix C.
@@
Table 2: Plots summarizing task performance (y-axis) and efficiency (x-axis) on Qwen3-4B model with various inference strategies. In all datasets, we find oracle strategy (i.e., the strategy that achieves the correct answer while incurring minimal tokens or the strategy incurring minimal tokens if no strategy leads to correct answer) significantly outperforms any uniform setting.
Inference Strategy Set We design the inference strategies to address the challenges revealed by our overthinking analysis. For Easy problems, we use lower temperature and reduced tokens to prevent the model from exploring unnecessary solution paths. Normal problems receive standard temperature with full token budget to enable comprehensive reasoning in the optimal region where the model’s uncertainty appropriately matches task complexity. For Hard problems, we implement a “Fail Fast” mechanism with strict token limits. Since our analysis indicates that these problems typically exceed the model’s capabilities regardless of generation length, this strategy prioritizes cutting computational losses on likely-to-fail queries to reallocate resources, rather than engaging in potentially unproductive reasoning. Table 2 summarizes the specific configurations for each reasoning strategy. Notably, these hyperparameters were empirically determined through a systematic grid search optimization rather than selected heuristically. Further details are provided in Appendix C.
Experimental Setting To establish the theoretical upper bound of our approach, we conduct an oracle experiment using Qwen3-4B across eight reasoning benchmarks. We evaluate on five mathematical reasoning tasks: GSM8K (Cobbe et al., 2021), MATH500 (Lightman et al., 2023), AIME 2024 & 2025, and OlympiadBench (He et al., 2024), along with three out-of-domain benchmarks: Minerva (Lewkowycz et al., 2022b), GPQA (Rein et al., 2024), and MMLU-Pro (Wang et al., 2024). For each input, we generates outputs from each of the three inference strategies, Easy, Normal, and Hard. The Oracle strategy is determined through a two-step process: first, we identify all strategies that yield a correct answer, and from this set, we then select the one with the minimal token consumption. The parameter configurations are set as the same as the fixed strategies in Table 2, and the max token limit is 32K. More visualization details can be found in Appendix G.
@@ 5.1 Stage 1: Data Generation with Proxy Model Sampling
For each problem $x$, we prompt the model to generate complete reasoning steps and final results with a maximum length of 32K tokens. We then compute the model’s uncertainty using the same generation entropy calculation described in Section 3. We perform this process with 10 sampling iterations per problem to ensure robust entropy estimates. This generation entropy serves as a proxy for task complexity. Grounded in the U-shaped entropy pattern in Section 3 and the Pareto-optimal Oracle analysis (Appendix D.3), we adopt the following heuristic labeling rule: Normal if correctness $\geq\,\alpha$ and entropy $\leq\,\beta$; Hard if correctness $\gamma$; Easy otherwise (typically the low-difficulty anomaly with moderate correctness but unexpectedly high uncertainty). The thresholds $\alpha,\beta,\gamma$ are set per model using a simple heuristic informed by the observed entropy–correctness distributions; we optionally perform a light sanity check on a small validation split to ensure label stability. Per-model values are reported in Appendix D.4.
@@ 6.1 Experimental Setup
Experimental Design. To demonstrate the probe’s ability to perform difficulty-adaptive reasoning, we mix benchmarks of different difficulty levels. We present averaged results across benchmarks with varying domains, in-domain (GSM8K, MATH500, AIME24&25, OlympiadBench) and out-of-domain (Minerva, GPQA, MMLU-Pro) to show adaptive performance across the difficulty spectrum. Each experiment is run three times; we report the mean across runs. For evaluation, we follow reliability protocols from Ye et al. (2025) [^2] https://github.com/GAIR-NLP/LIMO,Chen et al. (2025) [^3] https://github.com/IAAR-Shanghai/xVerify. Max token limits: for each model–benchmark pair, we first allow a generous cap (e.g., 32K) to observe the model’s longest output and then set a nearby rounded value as the per-benchmark max_tokens used in plots; see Appendix Table 13 for the finalized values and Appendix D.6 for details.
@@ 6.2 Reasoning LLMs Results
Performance Analysis DiffAdapt consistently outperforms fixed strategies across all model architectures and domains. On in-domain tasks (Figure 3(a)), Qwen3-4B shows the largest improvements, particularly at higher token budgets. DeepSeek-R1-Qwen-7B demonstrates stable gains, while DeepSeek-R1-Llama-8B exhibits significant benefits across model families. On out-of-domain evaluation (Figure 3(b)), the framework remains effective: DiffAdapt delivers consistent improvements across models and domains, with gains becoming more pronounced as the maximum token budget increases. By contrast, the training-free dynamic baseline DEER performs comparably to the strongest fixed strategy on in-domain datasets but still lags behind DiffAdapt; under distribution shift (Minerva, MMLU-Pro, GPQA), DEER exhibits limited generalization and larger performance degradation. These results confirm DiffAdapt’s robustness across different architectures, scales, and domains.
Key Findings The results reveal distinct performance patterns for different strategies. Easy strategies achieve high performance with minimal tokens but plateau quickly. Normal and Hard strategies improve continuously with increased computational budgets. This validates averaging across benchmarks of varying difficulty, as it captures the distinct computational requirements of different problem complexities. DiffAdapt exploits these patterns through adaptive strategy selection, consistently outperforming fixed approaches across all token limits. Compared with DEER, DiffAdapt delivers larger and more stable gains across token budgets and domains, indicating that difficulty-aware strategy selection generalizes better than confidence-based early exit. Models with larger inter-strategy performance differences show greater DiffAdapt improvements, confirming that adaptive selection benefits from strategy diversity.
@@ 6.4 Computational Efficiency Analysis
Token cost From Table 3, we observe substantial efficiency gains from DiffAdapt across models: Qwen3-4B achieves 22.4% token savings, while DS-R1-Qwen-7B and ThinkPrune-7B reduce usage by around 10%. In contrast, DEER increases token usage relative to Medium (e.g., 27.5-27.5% on Qwen3-4B and 53.3-53.3% on DS-R1-Qwen-7B). The reason is mechanistic: DEER operates under a fixed token budget and chooses continuation length primarily by confidence/probability, which frequently drives generations to the maximum token cap rather than allocating tokens adaptively by problem difficulty.
@@
7 Ablation Studies and Robustness Analysis
To rigorously validate the design choices and robustness of DiffAdapt, we conducted extensive ablation studies regarding hyperparameter sensitivity, probe architecture, and reasoning integrity.
7.1 Robustness and Design Choices
We investigate three key dimensions: (1) Threshold Sensitivity: whether the method requires manual tuning per model; (2) Probe Architecture: whether a non-linear MLP is necessary; and (3) Data Efficiency: performance impact of reducing training data. Our evaluation protocol aligns with Section 6, utilizing in-domain benchmarks including GSM8K, MATH500, AIME 24&25, and OlympiadBench.
Sensitivity to Thresholds ($\alpha,\beta,\gamma$) A key concern for deployment is whether the difficulty thresholds require fine-grained calibration. To test this, we replaced the optimized thresholds of Qwen3-4B with a configuration directly transferred from the DeepSeek-R1 model family ($\alpha=0.85,\beta=0.35,\gamma=0.60$). As shown in Table 5 (“Transferred Config”), the performance difference is negligible (avg. difference $\approx 0.3\%$). This confirms that DiffAdapt is highly robust to hyperparameter variations, enabling “plug-and-play” deployment without per-model re-calibration.
Probe Architecture and Data Scale Table 5 also highlights the impact of probe design. Replacing our 2-layer MLP with a simple Linear Head leads to a consistent accuracy drop ($\sim$3.2%), justifying the need for a lightweight non-linear classifier. Conversely, reducing the training data to only 30% results in minimal degradation, demonstrating that DiffAdapt is extremely data-efficient compared to RL-based methods that typically require large-scale rollout data.
Table 5: Ablation study on Qwen3-4B. We compare the Default DiffAdapt configuration against: (a) Transferred Thresholds from DeepSeek-R1 (to test robustness), (b) a Linear Probe (to test architecture necessity), and (c) 30% Training Data (to test data efficiency).
Baseline Robustness Check Probe Design Ablation
Token Budget DiffAdapt (Default) Transferred (DeepSeek-R1) Linear Head 30% Data
33.3% 59.3 59.9 56.4 64.5
50.0% 67.2 66.8 64.1 67.9
66.7% 72.4 72.7 68.7 69.3
83.3% 75.8 76.0 72.2 69.8
100% 76.8 76.6 74.0 70.1
Average 70.9 71.2 67.7 68.5
7.2 Reasoning Integrity Analysis
To address the concern that aggressive token reduction might compromise the logical completeness of reasoning chains (e.g., causing early truncation), we conducted a blind, pairwise LLM-as-a-Judge study, more details can be found in Appendix E.
Table 6: Pairwise comparison of reasoning quality (N=50). A blind Judge (Qwen3-30B) compared DiffAdapt vs. Baseline.
Outcome Count Percentage
DiffAdapt Wins 38 76%
Baseline Wins 6 12%
Tie 6 12%
Table 7: Failure analysis of the cases where Baseline won. “Truncation Error” indicates actual logic failure.
Reason for Loss Frequency
Subjective Preference 10% (5/50)
Truncation Error 2% (1/50)
We sampled 50 queries from GSM8K and used Qwen3-30B-A3B as an impartial judge to evaluate anonymized outputs from DiffAdapt and the Baseline (Normal strategy). The judge was explicitly instructed to penalize logical gaps. As shown in Table 6, DiffAdapt was preferred in 76% of cases, with the judge often citing that DiffAdapt produced “more concise and direct” reasoning without redundancy. Notably, catastrophic failure due to early truncation occurred in only 2% of cases (Table 7), refuting the concern that efficiency comes at the cost of reasoning integrity.
@@ 8 Conclusion
Reproducibility Statement
We aim for full reproducibility. Upon publication, we will release code, prompts, evaluation scripts, and configuration files to reproduce all tables and figures. We specify all random seeds, sampling settings (temperature/top-pp/number of samples n=10n{=}10), and the base token limit D.6. We will provide instructions for dataset preparation (e.g., DeepMath-103K splits and filtering), model versions used (Qwen3-4B, DeepSeek-R1-Qwen-7B, DeepSeek-R1-Llama-8B, Nemotron-1.5B, ThinkPrune-7B), and hardware/software environments needed to replicate results. All plots in the paper are generated by released scripts.
Ethics Statement
This work studies compute-efficient reasoning strategies for LLMs on public, decontaminated reasoning datasets. No personally identifiable or sensitive data are used. We discuss potential risks of misinterpretation and over-reliance on automatic reasoning systems and recommend careful human oversight in high-stakes scenarios. We will document evaluation limitations and known failure modes, and we avoid claims beyond the evaluated settings.
@@ Acknowledgements
We thank the members of NYU ML2 for their helpful feedback. This work was supported in part through the NYU IT High Performance Computing resources, services, and staff expertise. The work is partially funded by NSF CAREER award 2443271 and NSF award RI-2521091; the National Natural Science Foundation of China (Grant No. 62506318 and 62272122); the Guangdong Provincial Department of Education Project (Grant No. 2024KQNCX028); the Scientific Research Projects for the Higher-educational Institutions (Grant No. 2024312096) and the Guangzhou-HKUST(GZ) Joint Funding Program (Grant No. 2025A03J3957) from the Education Bureau of Guangzhou Municipality; and Hong Kong CRF grants under Grant No. C7004-22G and C6015-23G.
@@
Appendix A Use of Large Language Models
We used LLMs solely to aid and polish the writing (e.g., wording refinement and grammar), without generating or altering experimental designs, methods, results, or conclusions. All technical content, analyses, figures, and tables were authored and verified by the researchers.
Appendix B Limitations and Future Work
Transferability and Calibration. While our sensitivity analysis demonstrates that strategy thresholds are robust across model families (e.g., transferring parameters from DeepSeek-R1 to Qwen3 yields negligible $0.3\%$ deviation), extremely distinct domains may still benefit from a lightweight calibration phase. Future work could explore completely calibration-free mechanisms.
Prefill-only Trade-off. Our probe relies solely on prefill hidden states to predict difficulty. While this design minimizes inference latency by avoiding interruption of the decoding process, it inherently ignores generation-time dynamics that might reveal emerging complexity. A promising future direction is to incorporate lightweight generation signals (e.g., early step-wise entropy), though this introduces an efficiency-precision trade-off that must be carefully managed.
@@ Appendix B Limitations and Future Work
Appendix C Complete Reasoning Strategy
@@
C.1 Easy Strategy Prompt
@@
C.2 Normal Strategy Prompt
@@
C.3 Hard Strategy Prompt
For problems at the capability limit, we focus on efficient resource utilization and early termination of futile paths:
@@ C.3 Hard Strategy Prompt
  • Approach: Strategic method outline to implement a “Fail Fast” mechanism
C.4 Design Rationale
@@
  • Easy strategy discourages overthinking by emphasizing directness and verification rather than extensive exploration.
  • Normal strategy encourages systematic reasoning with full computational resources for optimal performance.
  • Hard strategy prioritizes resource conservation, identifying likely-to-fail queries early to avoid getting stuck in unproductive reasoning loops.
C.5 Hyperparameter Optimization
To determine the optimal configuration for these strategies, we avoided heuristic selection and instead conducted a comprehensive Grid Search experiment on the MATH500 dataset.
Search Space: We evaluated 125 distinct parameter combinations ($5_{\text{Normal}}\times 5_{\text{Hard}}\times 5_{\text{Easy}}$), varying Temperature ($T\in[0.1,1.2]$) and Max Token Ratios ($L\in[0.25\times,1.0\times]$).
Selection Criterion: We employed a constrained optimization approach:
  1. 1.
    Filter by Accuracy: We first identified all parameter combinations that maintained high accuracy ($\geq 95\%$) on the validation set.
  2. 2.
    Minimize Cost: From these candidates, we selected the configuration that yielded the lowest average token consumption (960.5 tokens).
Result: Table 8 demonstrates that our chosen configuration is empirically optimal, outperforming heuristic baselines in efficiency while preserving top-tier accuracy.
Table 8: Comparison of Top-Performing Strategy (Ours) vs. Heuristic Baselines on MATH500. Our grid-search tuned configuration achieves the best trade-off between accuracy and efficiency.
Strategy Configuration Description Acc (%) Avg Toks Max Toks Insight
[…]
Baseline A (Conservative) Heuristic uniform conservative (TT=0.6, full length for all) 94.0 1,003.2 12,461 Slightly lower accuracy; higher token cost.
Baseline B (Aggressive) Heuristic uniform high temp (TT=1.2, full length for all) 88.0 1,274.8 32,626 Suffers from “reasoning loops” on hard queries.
[…]